OpenStudy (anonymous):

Consider the line y=3x-3. Find the equation of the line that is parallel to this line and passes through the point (-8,6). Find the equation of the line that is perpendicular to this line and passes through the point (-8,6) Equation of parallel line: ? Equation of perpendicular line: ?

OpenStudy (anonymous):

I have no idea how to solve it?!

OpenStudy (anonymous):

i wish any buddy can help solve it

OpenStudy (marigirl):

Yes I can help you.

OpenStudy (marigirl):

Consider the line y=3x-3. Find the equation of the line that is parallel to this line and passes through the point (-8,6). ans: if a line is parallel, then that means both the lines have the same gradient. in this case y=mx+c .. the gradient is m. m=3. now take that information and plug it into this equation to create an equation y-y1=m(x-x1) where y1 is 6 m=3 x1=-8 y-6=3(x--8) y-6=3(x+8) Expand y-6=3x+24 plus 6 to both sides y=3x+30 therefore. 3x+30 is a line parallel to the line y=3x-3 going through (-8,6)

OpenStudy (anonymous):

thank you so much this is so helpful i understood it, so the parallel line is y=3x+30, how about the perpendicular line whats going to be the answer for it.

OpenStudy (marigirl):

Consider the line y=3x-3.Find the equation of the line that is perpendicular to this line and passes through the point (-8,6) ans: if a line is perpendicular, then multiplying both their gradients will result to -1 as mentioned before the gradient of y=3x-3 is m=3 m1 * m2= -1 where m1 is gradient of original line y=3x-3 and m2 is the gradient of the perpendicular line going thorough (-8,6) - which we dont know yet. but we can solve for it m1 * m2 = -1 3 * m2 = -1 m2 =-1/3 so now we have the gradient of the perpendicular line which is -1/3 so now plug this information into the equation y-y1=m(x-x1) to obtain our equation gradient= -1/3 x1= -8 y1=6 y-y1=m(x-x1) y-6=-1/3(x--8) y-6=-1/3(x+8) y-6=-1/3x-8/3 \[y=\frac{ -1 }{ 3 }x+\frac{ 10 }{ 3 }\]

OpenStudy (anonymous):

aha! now i got it.

OpenStudy (marigirl):

thanks! :) sorry i got carried away doing the maths (lol sometimes i get all excited doing the maths to get the answer) haha

OpenStudy (anonymous):

actually thanks to you for all what you have done. you made it simple to me